# Cross-correlating IceCube neutrinos with a large set of galaxy samples around redshift $z\sim 1$

### v1

@@ I Introduction
Cross-correlations with large scale structure (LSS) can also be used to search for neutrino sources [12, 13], assuming only that the neutrino sources reside in galaxies or galaxy clusters, which are known to be tracers of large scale structure. For a cross-correlation analysis, there is no need to design a source catalog that is likely to match actual neutrino sources, instead the method exploits the fact that matter is highly clustered on large scales in the Universe and thus any tracers of LSS (e.g. galaxies) are spatially correlated with each other. Cross-correlation analyses are ubiquitous in cosmology in terms of both the two-point correlation function in real space or the power spectrum in Fourier or harmonic space. Ref. [12] first proposed using the cross-correlation between neutrinos and galaxy samples to constrain the fraction of observed neutrinos that are correlated with LSS over a certain redshift range.
In this paper we build on the analysis of Ref. [12] by including several additional important components. We include modelling of the IceCube point spread function (PSF) and show that it has a very significant impact on the measured cross-correlation. We additionally model the details of galaxy clustering to illustrate how different galaxy samples can be sensitive to the redshift distribution of neutrino sources. We conduct an analysis of the 10-year IceCube public dataset of track-like events and provide forecasts for how well a future detector could constrain the properties of the neutrino source population.
@@ II.1 IceCube neutrinos
IceCube detects neutrinos through multiple different channels. Here we focus on “track-like” events, where a muon, produced by the interaction of a muon neutrino with the Antarctic ice, produces a long track of Cherenkov radiation that can be reaonably accurately traced back to an angular position on the sky.
The 10-year IceCube point-source data release [^1] http://doi.org/DOI:10.21234/sxvs-mt83 [19] consists of a catalog of just over a million track-like neutrino candidates detected between April 2008 and July 2018 with sky coordinates, reconstructed energies, and estimated angular uncertainties. The event selection was designed to identify tracks from high-energy muons passing through the IceCube detector and uses different selection criteria in the Northern and Southern skys.
@@ II.2 Galaxy samples
For our cross-correlation analysis, we look for large galaxy samples that span a variety of redshift ranges and have wide and uniform coverage in the northern sky. We provide details on the samples used below and their properties are summarized in Table 1.
@@ II.2.1 WISE-SuperCOSMOS
WISE-SuperCOSMOS [^2] http://ssa.roe.ac.uk/WISExSCOS.html [WI-SC, 20] is a full-sky low-redshift photometric galaxy catalog. These galaxies were obtained by cross-matching data from the Wide-field Infrared Survey Explorer [WISE, 21] with the photographic plates of SuperCOSMOS [22, 23]. WI-SC contains $\sim 20$ million sources up to redshift of $z=0.4$. The photometric redshifts have a mean error of $\sigma_{z}/(1+z)\approx 0.035$. We divide the catalog into 5 equal-size redshift bins between $z=0.1$ and $z=0.35$ (labelled as WI-SC 1 through 5 in Table 1). We estimate the redshift distribution of each bin by convolving the photometric redshift distributions with Gaussians given by the mean photometric redshift errors. We use the sky mask constructed by Ref. [20] that masks the galactic disk, bulge, the Magellanic clouds, and other areas of large stellar contamination or dust extinction. As noted in Refs. [24, 25], some stellar contamination and obscuration still remains at the few per cent level, but we treat this as negligible due to the fact that our measurements are severely limited by the quality of the neutrino data.
@@ II.2.2 DESI LRGs
To probe redshifts between $z\sim 0.4$ and $z\sim 1$, we use luminous red galaxy (LRG) samples selected from DESI Legacy Imaging Surveys DR9333https://www.legacysurvey.org/dr9/description/ by Ref. [26]. These samples have wide sky coverage and spectroscopic redshift distributions, and are designed for cross-correlation applications with 4 tomographic bins. We specifically use the “extended LRG” sample that has 2-3 times as many galaxies as the main sample presented in Ref. [26] in order to gain lower shot noise at the expense of a slighlty less clean sample. In Table 1, these samples are labelled as DESI 1–4.
@@ II.2.5 CIB maps
In addition to galaxy catalogs, we use the large-scale CIB maps generated by Ref. [34] based on the Planck High-Frequency Instrument sky maps at 353, 545, and 857 GHz. Ref. [34] removed the Galactic dust emission from these maps using a template based on H I column desnity. These maps trace the integrated extragalactic infrared emission and thus act as maps of unresolved galaxies over a very wide redshift range.
@@ III Theory
We are interested in modelling the angular cross-spectrum between galaxies with known properties and some unknown astrophysical neutrino sources. In both cases, the projected overdensity field on the sky $\delta(\hat{\mathbf{n}})$ is related to the 3D overdensity field $\Delta(\mathbf{x},t)$
@@
In the case of the neutrinos, the situation is complicated by the very large background of atmospheric neutrinos. Assuming that the atmospheric neutrinos are uniformly distributed across the sky and uncorrelated with the astrophysical ones, $b_{\nu}$ is completely degenerate with $f_{\text{astro}}$, the fraction of neutrinos that are astrophysical [see appendix A in 12]. This is the exact same effect that an unclustered stellar contamination has on the measurement of galaxy bias [27]. Since astophysical and atmospheric neutrinos are completely indistinguishable on an event-by-event basis, the combination $b_{\nu}f_{\text{astro}}$ determines the amplitude of $C_{\ell}^{g\nu}$.
@@
The neutrino auto-spectrum is modelled as
@@
The galaxy-neutrino cross-spectrum is modelled as
@@
We calculate all of these theory quantities using the Core Cosmology Library [^8] https://github.com/LSSTDESC/CCL (CCL) [38]. The matter power spectrum $P_{m}(k,z)$ is calculated using the HaloFit model [39] as implemented in CCL. The remaining components that go into the theory prediction for the galaxy-neutrino cross-spectrum (Eq. 6) for a given galaxy sample are: the galaxy bias $b_{g}$ and redshift distrribution $p_{g}(z)$, the neutrino source redshift distribution $p_{\nu}(z)$, and the neutrino clustering amplitude $b_{\nu}f_{\text{astro}}$.
@@
The CIB maps have to be treated slightly differently since they do not consist of resolved galaxies. Ultimately the CIB model still consists of a galaxy bias and redshift distribution, but the redshift distribution is written in terms of the infrared emissivity of galaxies across redshift and the cosmic star-formation rate desnity. We directly follow Section 2 of Ref. [35] to construct a linear model of the CIB anisotropies and we use the best fit parameters reported in their Table 3.
@@ IV.1 Maps
For the neutrinos, as a first step, we consider all events with reconstructed energies $E\geq 1$ TeV. This selection removes the lowest energy events that are expected to be dominated by atmosperic neutrinos while keeping a fairly large sample ($\sim 400,000$ events) to minimize statistical errors. Later, we also present results for neutrinos binned by energy, but the analysis steps are exactly the same as described here for the single bin. We construct a smoothed neutrino counts map $N_{\nu}$ for a given event selection by treating each event as a 2D circular Gaussian centered at the reported angular position and using the reported angular uncertainty as the standard deviation. This procedure is essentially the same as used by the IceCube collaboration to construct spatial likelihoods for point source searches [44, 45]. In our analysis, constructing a smoothed counts map allows us to maximize the signal-to-noise ratio (SNR) by weighting events according to the quality of their localization.
Next, we use randomized data to construct an estimate of the local mean neutrino counts $\bar{N}_{\nu}$. At a given energy, the effective area of the IceCube detector depends only on the source declination (due to the unique position at the South Pole); we can generate random realizations of the neutrino count maps by randomizing the R.A. values of each event. We construct the mean counts map by averaging 50 randomized smoothed counts maps and then making the resuting map independent of R.A. by averaging over declination bands.
@@ IV.2 IceCube effective beam
While galaxy positions are generally measured at a precision of roughly a few arc-seconds, IceCube’s degree-scale sky localization of track-like events introduces very significant smoothing that must be taken into account by modelling the detector point-spread function (PSF, or beam window function in harmonic space). In general, the IceCube beam window function is neither constant nor isotropic, since the estimated event angular uncertainties depend on energy and declination.
@@
We estimate the effective beam directly from the data. Since our smoothed neutrino overdensity map effectively contains two powers of the PSF (one from the detector and one from the smoothing operation during map construction), we use
$B_{\ell}^{\text{eff}}=C_{\ell}^{\tilde{\nu}\tilde{\nu}}/C_{\ell}^{\nu\nu},$ (10)
where $\tilde{\nu}$ and $\nu$ represent the smoothed and non-smoothed neutrino maps respectively. In the simplest case, where the PSF does not vary significantly across the sky, this estimate should be equivalent to averaging the squares of the Gaussian beams calculated from each event’s angular uncertainty (the beam for each event is squared to account for the two powers of the PSF, as mentioned above). We show these two estimates of the effective beam in Fig. 3, where we additionally compare them to a Gaussian beam calculated from the median angular uncertainty of all selected events. We find that the two estimates of the effective beam agree very well, except at very large spatial scales where spatial variations in the beam become more significant. We also see that the effective beam is very non-Gaussian, even though we assumed that the PSF for each event was Gaussian. Correctly modelling this beam is important for any cross-correlation analyses.
To fully validate this estimate of the effective beam, we would need to do detailed simulations of the IceCube detector, but we have validated this procedure on simple realisations of Gaussian random fields. There is an inherent uncertainty in this estimate of the effective beam due to the assumption that all events can be described by circular Gaussians. As can be seen from the catalog of high-energy neutrino alerts [46], the more detailed event spatial likelihoods are very often non-circular and often non-Gaussian.
Another caveat is that Eqs. 8 and 9 are not necessarily both true. In the case of a strongly spatially varying beam, there could be significant differences between the effective beams defined in terms of cross-spectra with different galaxy samples. In our analysis, all the galaxy samples have fairly similar coverage in the Northern sky and we have explicitly checked that the result of Eq. 10 does not significantly change when computing the power spectra using the overlap of the neutrino mask and the different galaxy sample masks.
@@
Figure 3: Effective beam for the smoothed neutrino map consisting of all events with $E>1$ TeV. The analytic estimate is calculated by averaging the Gaussian beams of each event. We also plot a Gaussian beam corresponding to the median angular uncertaintiy of all selected events to show that the overall effective beam is highly non-Gaussian. Essentially no information remains in the power spectra past $\ell\sim 350$.
IV.3 Angular power spectra and covariance matricies

### v2

@@ I Introduction
Cross-correlations with large scale structure (LSS) can also be used to search for neutrino sources [12, 13], assuming only that the neutrino sources reside in galaxies or galaxy clusters, which are known to be tracers of large scale structure. For a cross-correlation analysis, there is no need to design a source catalog that is likely to match actual neutrino sources, instead the method exploits the fact that matter is highly clustered on large scales in the Universe and thus any tracers of LSS (e.g. galaxies) are spatially correlated with each other. Cross-correlation analyses are ubiquitous in cosmology in terms of both the two-point correlation function in real space or the power spectrum in Fourier or harmonic space. Ref. [12] first proposed using the cross-correlation between neutrinos and galaxy samples to constrain the fraction of observed neutrinos that are correlated with LSS over a certain redshift range, while Ref. [13] conducted a detailed analysis of the cross-correlation between neutrinos and the gamma-ray background including a quantification of the IceCube point spread function (PSF).
In this paper we build on the analysis of Ref. [12] by including several additional important components. Similar to Ref [13], we include modeling of the IceCube PSF and show that it has a very significant impact on the measured cross-correlation. But, we use additional event-level smoothing in order to maximize the potential signal-to-noise and attempt to account for the spatially varying PSF. We additionally model the details of galaxy clustering to illustrate how different galaxy samples can be sensitive to the redshift distribution of neutrino sources. We conduct an analysis of the 10-year IceCube public dataset of track-like events and provide forecasts for how well a future detector could constrain the properties of the neutrino source population.
@@ II.1 IceCube neutrinos
IceCube detects neutrinos through multiple different channels. Here we focus on “track-like” events, where a muon, produced by the interaction of a muon neutrino with the Antarctic ice, produces a long track of Cherenkov radiation that can be reasonably accurately traced back to an angular position on the sky.
The 10-year IceCube point-source data release [^1] http://doi.org/DOI:10.21234/sxvs-mt83 [19] consists of a catalog of just over a million track-like neutrino candidates detected between April 2008 and July 2018 with sky coordinates, reconstructed energies, and estimated angular uncertainties. The event selection was designed to identify tracks from high-energy muons passing through the IceCube detector and uses different selection criteria in the Northern and Southern skies.
@@ II.2 Galaxy samples
For our cross-correlation analysis, we look for large galaxy samples that span a variety of redshift ranges and have wide and uniform coverage in the northern sky. We provide details on the samples used below and summarize their properties in Fig. 1 and Table 1.
@@ II.2.1 WISE-SuperCOSMOS
WISE-SuperCOSMOS [^2] http://ssa.roe.ac.uk/WISExSCOS.html [WI-SC, 20] is a full-sky low-redshift photometric galaxy catalog. These galaxies were obtained by cross-matching data from the Wide-field Infrared Survey Explorer [WISE, 21] with the photographic plates of SuperCOSMOS [22, 23]. WI-SC contains $\sim 20$ million sources up to redshift of $z=0.4$. The photometric redshifts have a mean error of $\sigma_{z}/(1+z)\approx 0.035$. We divide the catalog into 5 equal-size redshift bins between $z=0.1$ and $z=0.35$ (labeled as WI-SC 1 through 5 in Table 1). We estimate the redshift distribution of each bin by convolving the photometric redshift distributions with Gaussians given by the mean photometric redshift errors. We use the sky mask constructed by Ref. [20] that masks the galactic disk, bulge, the Magellanic clouds, and other areas of large stellar contamination or dust extinction. As noted in Refs. [24, 25], some stellar contamination and obscuration still remains at the few per cent level, but we treat this as negligible due to the fact that our measurements are severely limited by the quality of the neutrino data.
@@ II.2.2 DESI LRGs
To probe redshifts between $z\sim 0.4$ and $z\sim 1$, we use luminous red galaxy (LRG) samples selected from DESI Legacy Imaging Surveys DR9333https://www.legacysurvey.org/dr9/description/ by Ref. [26]. These samples have wide sky coverage and spectroscopic redshift distributions, and are designed for cross-correlation applications with 4 tomographic bins. We specifically use the “extended LRG” sample that has 2-3 times as many galaxies as the main sample presented in Ref. [26] in order to gain lower shot noise at the expense of a slightly less clean sample. In Table 1, these samples are labeled as DESI 1–4.
@@ II.2.5 CIB maps
In addition to galaxy catalogs, we use the large-scale CIB maps generated by Ref. [34] based on the Planck High-Frequency Instrument sky maps at 353, 545, and 857 GHz. Ref. [34] removed the Galactic dust emission from these maps using a template based on H I column density. These maps trace the integrated extragalactic infrared emission and thus act as maps of unresolved galaxies over a very wide redshift range.
@@
Figure 1: Redshift distributions for all of the galaxy samples used in this analysis. A summary of the properties of each galaxy sample is given in Table 1.
@@ III Theory
We are interested in modeling the angular cross-spectrum between galaxies with known properties and some unknown astrophysical neutrino sources. In both cases, the projected overdensity field on the sky $\delta(\hat{\mathbf{n}})$ is related to the 3D overdensity field $\Delta(\mathbf{x},t)$
@@
In the case of the neutrinos, the situation is complicated by the very large background of atmospheric neutrinos. Assuming that the atmospheric neutrinos are uniformly distributed across the sky and uncorrelated with the astrophysical ones, $b_{\nu}$ is completely degenerate with $f_{\text{astro}}$, the fraction of neutrinos that are astrophysical [see appendix A in 12]. This is the exact same effect that an unclustered stellar contamination has on the measurement of galaxy bias [27]. Since astrophysical and atmospheric neutrinos are completely indistinguishable on an event-by-event basis, the combination $b_{\nu}f_{\text{astro}}$ determines the amplitude of $C_{\ell}^{g\nu}$.
@@
The neutrino auto-spectrum is modeled as
@@
The galaxy-neutrino cross-spectrum is modeled as
@@
We calculate all of these theory quantities using the Core Cosmology Library [^8] https://github.com/LSSTDESC/CCL (CCL) [38]. The matter power spectrum $P_{m}(k,z)$ is calculated using the HaloFit model [39] as implemented in CCL. The remaining components that go into the theory prediction for the galaxy-neutrino cross-spectrum (Eq. 6) for a given galaxy sample are: the galaxy bias $b_{g}$ and redshift distribution $p_{g}(z)$, the neutrino source redshift distribution $p_{\nu}(z)$, and the neutrino clustering amplitude $b_{\nu}f_{\text{astro}}$.
@@
The CIB maps have to be treated slightly differently since they do not consist of resolved galaxies. Ultimately the CIB model still consists of a galaxy bias and redshift distribution, but the redshift distribution is written in terms of the infrared emissivity of galaxies across redshift and the cosmic star-formation rate density. We directly follow Section 2 of Ref. [35] to construct a linear model of the CIB anisotropies and we use the best fit parameters reported in their Table 3.
@@ IV.1 Maps
For the neutrinos, as a first step, we consider all events with reconstructed energies $E\geq 1$ TeV. This selection removes the lowest energy events that are expected to be dominated by atmospheric neutrinos while keeping a fairly large sample ($\sim 400,000$ events) to minimize statistical errors. Later, we also present results for neutrinos binned by energy, but the analysis steps are exactly the same as described here for the single bin.
We construct a smoothed neutrino counts map $N_{\nu}$ for a given event selection by treating each individual event as a 2D circular Gaussian centered at the reported angular position and using the reported angular uncertainty as the standard deviation. This procedure is similar to that used by the IceCube collaboration to construct spatial likelihoods for point source searches [44, 45]. In our analysis, constructing a smoothed counts map allows us to maximize the signal-to-noise ratio (SNR) by weighting events according to the quality of their localization in a manner very similar to a Wiener filter. We have confirmed that our final results are consistent whether we use the smoothed map described here or the raw un-smoothed neutrino counts map, but using the smoothed map results in tighter final constraints showing that the smoothing operation results in a closer-to-optimal analysis.
Next, we use randomized data to construct an estimate of the local mean neutrino counts $\bar{N}_{\nu}$. At a given energy, the effective area of the IceCube detector depends only on the source declination (due to the unique position at the South Pole); we can generate random realizations of the neutrino count maps by randomizing the R.A. values of each event. We construct the mean counts map by averaging 50 randomized smoothed counts maps and then making the resulting map independent of R.A. by averaging over declination bands.
@@ IV.2 IceCube effective beam
While galaxy positions are generally measured at a precision of arc-seconds, IceCube’s degree-scale sky localization of track-like events introduces very significant smoothing that must be taken into account by modeling the detector point-spread function (PSF, or beam window function in harmonic space). In general, the IceCube beam window function is neither constant nor isotropic, since the estimated event angular uncertainties depend on energy and declination.
@@
Under the assumption that the PSF of each individual event is a circular Gaussian and that it only varies on very large scales across the sky, we can estimate the effective beam directly from the data. The method we use is as follows. First, we assume that we can write the neutrino map as observed by IceCube (i.e., before our additional smoothing) as a sum over $N$ sets of events, indexed by $i$, that share the same PSF:
$\nu=\frac{1}{N}\sum_{i}^{N}(B^{i}\otimes s+n^{i}),$ (10)
where $B^{i}$ is the PSF of event set $i$, $\otimes$ represents the convolution operation, $s$ is some common signal map representing the clustering of astrophysical neutrinos, and $n^{i}$ is the non-clustered noise of event set $i$ representing shot noise and/or the atmospheric foreground. After our smoothing operation which smooths each event with the reported angular uncertainty, the neutrino map gains a second power of the PSF:
$\tilde{\nu}=\frac{1}{N}\sum_{i}B^{i}\otimes(B^{i}\otimes s+n^{i}).$ (11)
In harmonic space, the convolutions become products, so we can write
$\nu_{\ell m}$ $=\frac{1}{N}\sum_{i}(B^{i}_{\ell}s_{\ell m}+n^{i}_{\ell m})$ (12)
$\tilde{\nu}_{\ell m}$ $=\frac{1}{N}\sum_{i}\left((B^{i}_{\ell})^{2}s_{\ell m}+B^{i}_{\ell}n^{i}_{\ell m}\right).$ (13)
Then, the auto-spectrum of the map observed by IceCube is given by
$\langle\nu_{\ell m}^{2}\rangle=\overline{B_{\ell}}^{2}\langle s_{\ell m}^{2}\rangle+\langle\overline{n_{\ell m}}^{2}\rangle,$ (14)
where we have used the bar to denote the average over the $N$ sets of neutrino events.
Similarly, the auto-spectrum of the neutrino map that has been smoothed at the event level is given by
$\langle\tilde{\nu}_{\ell m}^{2}\rangle=\overline{B_{\ell}^{2}}^{2}\langle s_{\ell m}\rangle+\langle\overline{B_{\ell}n_{\ell m}}^{2}\rangle.$ (15)
The second term in the above equation can be simplified since noise between events are uncorrelated, or $\langle n^{i}_{\ell m}n^{j}_{\ell m}\rangle=0$ for $i\neq j$:
$\langle\overline{B_{\ell}n_{\ell m}}^{2}\rangle$ $=\langle\frac{1}{N}\sum_{i}\left(B^{i}_{\ell}n^{i}_{\ell m}\right)^{2}+\text{cross terms}\rangle$ (16)
$=\overline{B_{\ell}^{2}}\langle\overline{n^{2}_{\ell m}}\rangle.$ (17)
The same trick can be used on the second term in Eq 14, so we get
$\langle\nu_{\ell m}^{2}\rangle$ $=\overline{B_{\ell}}^{2}\langle s_{\ell m}^{2}\rangle+\langle\overline{n_{\ell m}^{2}}\rangle$ (18)
$\langle\tilde{\nu}_{\ell m}^{2}\rangle$ $=\overline{B_{\ell}^{2}}^{2}\langle s_{\ell m}\rangle+\overline{B_{\ell}^{2}}\langle\overline{n_{\ell m}^{2}}\rangle.$ (19)
Finally, if we have some other sample $g$ (for example, some galaxy sample) that also traces the same signal map $s$, the cross-spectra with the neutrino maps are
$\langle\nu_{\ell m}g_{\ell m}\rangle$ $=\overline{B_{\ell}}\langle s_{\ell m}^{2}\rangle$ (20)
$\langle\tilde{\nu}_{\ell m}g_{\ell m}\rangle$ $=\overline{B_{\ell}^{2}}\langle s_{\ell m}\rangle.$ (21)
Comparing Eqs. 18-21, we come up with two methods to estimate the effective beam $B_{\ell}^{\text{eff}}$. First, from Eqs. 19 and 21 we see that the effective beam of our smoothed neutrino map can be estimated simply by averaging the squares of the individual Gaussian beams of each event:
$B_{\ell}^{\text{eff}}=\overline{B_{\ell}^{2}}.$ (22)
We call this the analytic estimate of the effective beam.
Secondly, by comparing Eqs. 18 and 19, we can see that the effective beam can also be estimated by looking at the ratio of the auto-spectra of the two neutrino maps. Since we are very much in the noise dominated regime, $\langle s_{\ell m}^{2}\rangle$ is small relative to the white noise, so
$B_{\ell}^{\text{eff}}=\overline{B_{\ell}^{2}}=C_{\ell}^{\tilde{\nu}\tilde{\nu}}/C_{\ell}^{\nu\nu}.$ (23)
We treat this second estimate as our fiducial estimate of the effective beam, since we suspect that the auto-spectra might capture some of the information about spatial variations of the beam.
@@
Figure 4: Effective beam (accounting for the instrument PSF and our applied smoothing) for the smoothed neutrino map consisting of all events with $E>1$ TeV. The analytic estimate is calculated by averaging the Gaussian beams of each event. We also plot a Gaussian beam corresponding to the median angular uncertainty of all selected events to show that the overall effective beam is highly non-Gaussian. Essentially no information remains in the power spectra past $\ell\sim 350$.
We plot both estimates of the effective beam in Fig. 4 and find that they agree very well, except at very large spatial scales where spatial variations in the beam become more significant. We also see that the effective beam is very non-Gaussian, even though we assumed that the PSF for each event was Gaussian. Correctly modelling this beam is important for any cross-correlation analyses, see also Fig. 7 in Appendix A of Ref. [13] where the same conclusion is reached. While the effective beam estimated by Ref. [13] is significantly different from ours due to different event selections, when attempting to match their selection, we find that our final results shift by at most $0.4\sigma$ when using their estimated beam instead of ours. As a final note, if the raw neutrino map as observed by IceCube is used (as in Refs. [12, 13]) instead of our smoothed version, the effective beam should be estimated using
$B_{\ell}^{\text{eff}}=\overline{B_{\ell}}=C_{\ell}^{\tilde{\nu}\nu}/C_{\ell}^{\nu\nu},$ (24)
since there is only one power of the PSF present in the map instead of two as in our analysis.
There are several caveats to the above analysis, the rigorous treatment of which we leave for future work. First, while we have validated our estimator of the effective beam on simple Gaussian simulations, to fully validate it we would need detailed simulations of the IceCube detector which are not available publicly. There is an inherent uncertainty in this estimate of the effective beam due to the assumption that all events can be described by circular Gaussians. As can be seen from the catalog of high-energy neutrino alerts [46], the more detailed event spatial likelihoods are very often non-circular and often non-Gaussian. These effects will become more important in future analyses with better sensitivity. Additionally, in the case of a strongly spatially varying beam, there could be significant differences between the effective beams defined in terms of cross-spectra with different galaxy samples. In our analysis, all the galaxy samples have fairly similar coverage in the Northern sky and we have explicitly checked that the result of Eq. 23 does not significantly change when computing the power spectra using the overlap of the neutrino mask and the different galaxy sample masks.
IV.3 Angular power spectra and covariance matrices